# Noah Porcelli

### v1

@@ 1 Introduction
Let $(M,\omega)$ be a symplectic manifold, and $L\subseteq M$ a Lagrangian submanifold. In [13], Yau studied the following question:
@@
What diffeomorphisms $\tilde{\phi}$ of $L$ can be extended to a Hamiltonian diffeomorphism $\phi$ on $M$?
This question has been studied in various places. For example, some partial classification results in the case $L$ is a monotone 2-torus have been obtained in [13, 1]; in particular, even in simple cases $\tilde{\phi}$ can act non-trivially on $H^{*}(L)$. A more comprehensive history of this question can be found in the introduction in [1].
We now assume (for the rest of the paper) that $M$ is compact or Liouville, and that $L\subseteq M$ is a compact, relatively exact Lagrangian. We also assume all Hamiltonians are compactly supported: this is without any loss of generality, as can be seen by using an appropriate cut-off function supported away from the image of $L$ under the flow of this Hamiltonian.
@@
Let $\tilde{\phi}\mathrel{\mathop{:}}L\to L$ be a diffeomorphism, and assume that there is a Hamiltonian diffeomorphism $\phi\mathrel{\mathop{:}}M\to M$ extending $\tilde{\phi}$.
@@
    ([7]): $\phi$ acts as the identity on $H^{*}(L;\mathbb{Z}/2)$.
@@
    ([7, 10]): If $L$ is Spin, $\phi$ acts as the identity on $H^{*}(L;\mathbb{Z})$.
@@
    ([12]): If $L$ is oriented, $\phi$ is orientation-preserving.
Further results regarding the action of $\phi$ on string topology as well as some generalised cohomology theories were obtained in [10]. Motivated by these, one can conjecture:
@@
Consider the space $\mathcal{G}$ of Hamiltonian diffeomorphisms $\phi\mathrel{\mathop{:}}M\to M$ sending $L$ to itself. Then $\mathcal{G}$ is contractible.
The case where $L$ is the zero section in $M=T^{*}L$ would follow from the strong nearby Lagrangian conjecture.
1.1 Extension to the $C^{0}$ setting
@@
A Hamiltonian homeomorphism is a homeomorphism $\theta\mathrel{\mathop{:}}M\to M$ which is a $C^{0}$ limit of Hamiltonian diffeomorphisms.
It is known due to work of Eliashberg-Gromov [3, 4] that any Hamiltonian homeomorphism which is also a diffeomorphism, is in fact a Hamiltonian diffeomorphism. In general however the geometry of Hamiltonian homeomorphisms is much less understood than those of Hamiltonian diffeomorphisms.
The dynamical behaviour of Hamiltonian homeomorphisms is very different to Hamiltonian diffeomorphisms: for example, the Arnold conjecture fails (badly) for Hamiltonian homeomorphisms [2].
While Question 1.1 has not been studied in the $C^{0}$ setting, a similar question has been studied very recently in [8].
Here we extend some of the results of Theorem 1.2 to the $C^{0}$ setting (it would be interesting to extend the other results to the $C^{0}$ setting too).
@@
Let $\tilde{\phi}\mathrel{\mathop{:}}L\to L$ be a homeomorphism, and assume that there is a Hamiltonian homeomorphism $\phi\mathrel{\mathop{:}}M\to M$ extending $\tilde{\phi}$.
@@
Theorem 1.5 holds for various other generalised cohomology theories $R^{*}$ (such as real $K$-theory $KO^{*}$) provided the condition [10, Proposition 1.14] holds for $R^{*}$. Examples of diffeomorphisms of closed manifolds which act as the identity but not on (real or complex) $K$-theory are given in [10, Appendices A & B].
@@
If $M$ is Liouville and $L$ is exact, it would follow from the nearby Lagrangian conjecture that if $\tilde{\phi}$ is as in Theorem 1.5, then $\tilde{\phi}$ is homotopic to a diffeomorphism $L\to L$ which extends to a Hamiltonian diffeomorphism of $M$; Theorem 1.5 would then follow from Theorem 1.2.
@@ 1.2 Extension to the relative setting
Assume $\tilde{\phi}\mathrel{\mathop{:}}L\to L$ is a homeomorphism which extends to a Hamiltonian homeomorphism $\phi\mathrel{\mathop{:}}M\to M$. $\phi$ induces a map of pairs $(M,L)\to(M,L)$, and hence a map on relative cohomology. Even though $\phi$ is homotopic to the identity on $M$ (just as a map, not necessarily of pairs), and $\tilde{\phi}$ acts as the identity on cohomology, it does not immediately imply the map induced on relative cohomology is the identity, though Conjecture 1.3 would imply this. Nevertheless:
@@
Let $\tilde{\phi}\mathrel{\mathop{:}}L\to L$ be a homeomorphism, and assume that there is a Hamiltonian homeomorphism $\phi\mathrel{\mathop{:}}M\to M$ extending $\tilde{\phi}$.
@@
The proof is similar to the proof of Theorem 1.5, but whereas the proof of Theorem 1.5 uses a moduli space of holomorphic curves with domain $Z_{C}$, $\mathcal{P}_{C}$, along with the evaluation map on the boundary of $Z_{C}$, $ev\mathrel{\mathop{:}}\mathcal{P}\times\partial Z_{C}\to M$, the proof of Theorem 1.7 uses the evaluation map on the whole of $Z_{C}$, $ev\mathrel{\mathop{:}}\mathcal{P}\times Z_{C}\to M$.
@@ 1.3 Acknowledgements
The author thanks Amanda Hirschi, Ivan Smith and Yuhan Sun for helpful conversations, and Jack Smith for helpful conversation and useful comments on an earlier draft. The author is supported by the Engineering and Physical Sciences Research Council [EP/W015889/1].
@@ 2 Smooth approximation
We first choose a complete Riemannian metric on $M$, and let $03\varepsilon>$ 0 be its injectivity radius. Write $d$ for the Riemannian distance function.
Now choose a smooth Hamiltonian diffeomorphism $\psi^{1}\mathrel{\mathop{:}}M\to M$ which is sufficiently $C^{0}$-close to $\phi$ that for all $x$ in $M$, $d(\psi^{1}(x),\phi(x))\leq\varepsilon$. We extend $\psi^{1}$ to a 1-parameter family $\{\psi^{t}\}_{t\in[0,1]}$ of Hamiltonian diffeomorphisms such that $\psi^{0}$ is the identity. We write $L^{\prime}$ for $\psi^{1}(L)$.
Let $U$ be a tubular neighbourhood of $L$ and $\pi\mathrel{\mathop{:}}U\to L$ the projection map. We choose $U$ small enough that for all $x$ in $U$,
@@
We may choose $\psi^{1}$ $C^{0}$-close enough to $\phi$ such that $L^{\prime}$ lives in $U$, and generically so that $L$ and $L^{\prime}$ intersect transversally.
@@
The maps $\phi,\pi\circ\psi^{1}\mathrel{\mathop{:}}L\to L$ are homotopic.
@@
For $x$ in $L$, since $\pi\circ\phi(x)=\phi(x)$, by the triangle inequality $d(\phi(x),\pi\circ\psi^{1}(x))\leq 3\varepsilon$. Thus there is a unique small geodesic between $\phi(x)$ and $\pi\circ\psi^{1}(x)$; moving along this geodesic for each $x$ gives the required homotopy. ∎

### v2

@@ 1 Introduction
1.1 Background
Recall a symplectic manifold $(M,\omega)$ is a manifold $M$, equipped with a 2-form $\omega\in\Omega^{2}(M)$ which is closed ($d\omega=0$) and non-degenerate (meaning $\omega^{\mathrm{Dim}(M)/2}$ is everywhere non-zero, in particular $\mathrm{Dim}(M)$ must be even). Important examples include cotangent bundles $T^{*}Q$ and smooth complex quasiprojective varieties. Two of the main objects of study in classical symplectic geometry are Lagrangian submanifolds and Hamiltonian diffeomorphisms along with their topological and dynamical aspects.
A Lagrangian submanifold of a symplectic manifold $L\subseteq(M,\omega)$ is a submanifold satisfying $\omega|_{L}=0$, and $\mathrm{Dim}(L)=\mathrm{Dim}(M)/2$. Examples include the zero section $Q_{0}$ in a cotangent bundle $T^{*}Q$, the real loci of complex quasiprojective varieties defined by real polynomials, and the fibres of moment maps of smooth toric varieties.
A Hamiltonian diffeomorphism $\phi^{1}\mathrel{\mathop{\ordinarycolon}}M\to M$ is a diffeomorphism of $M$ which is the time-one flow of a Hamiltonian vector field $\{X^{t}_{H}\}_{t\in[0,1]}$, which is defined via the equation $\omega(\cdot,X^{t}_{H})=dH^{t}$ for some smooth map $H\mathrel{\mathop{\ordinarycolon}}[0,1]\times M\to\mathbb{R}$ (nondegeneracy of $\omega$ implies this is well-defined). These arise naturally in classical dynamics as a formulation of Hamiltonian’s equations of motion to abstract manifolds. We write $\operatorname{Ham}(M)$ for the space of Hamiltonian diffeomorphisms of $M$. We refer to [16] for more background on symplectic geometry.
In [24], Yau studied the following question on the topology of Lagrangian submanifolds:
@@ 1.1 Background
Let $L\subseteq(M,\omega)$ be a Lagrangian in a symplectic manifold. What diffeomorphisms $\tilde{\phi}$ of $L$ can be extended to a Hamiltonian diffeomorphism $\phi$ on $M$?
This question has been studied in various places. For example, some partial classification results in the case $L$ is a monotone Lagrangian 2-torus have been obtained in [24, 1]; in particular, even in simple cases $\tilde{\phi}$ can act non-trivially on $H^{*}(L)$. A more comprehensive history of this question can be found in the introduction in [1].
A Lagrangian $L$ is called relatively exact if the symplectic area of any disc with bounday on $L$ is 0: in symbols, $\omega\cdot\pi_{2}(M,L)=0$. Such Lagrangians enjoy much stronger rigidity properties than arbitrary Lagrangians; for example, in many cases they cannot be displaced from themselves by a Hamiltonian isotopy [5].
Assumption 1.2.
We now fix (for the rest of the paper) a symplectic manifold $(M,\omega)$ and a Lagrangian submanifold $L\subseteq M$.
We assume everywhere that $M$ is compact or Liouville (cf. [22, Section 7(b)]) [^2] Roughly, this says that $M$ behaves nicely outside a compact set.. We also assume that $L\subseteq M$ is a compact, relatively exact Lagrangian.
We also assume for simplicity that all Hamiltonians are compactly supported [^3] This is without any loss of generality, as can be seen by using an appropriate cut-off function supported away from the image of $L$ under the flow of this Hamiltonian..
@@
Let $L\subseteq M$ be a compact, relatively exact Lagrangian in a (compact or Liouville) symplectic manifold $M$.
Let $\tilde{\phi}\mathrel{\mathop{\ordinarycolon}}L\to L$ be a diffeomorphism, and assume that there is a Hamiltonian diffeomorphism $\phi\mathrel{\mathop{\ordinarycolon}}M\to M$ extending $\tilde{\phi}$.
@@
    [10]: $\phi$ acts as the identity on $H^{*}(L;\mathbb{Z}/2)$.
@@
    [10, 17]: If $L$ is Spin, $\phi$ acts as the identity on $H^{*}(L;\mathbb{Z})$.
@@
    [23]: If $L$ is oriented, $\phi$ is orientation-preserving.
  4. 4.
    [17]: $\phi$ acts as the identity on the set of conjugacy classes of $\pi_{1}L$.
Further results regarding the action of $\phi$ on string topology as well as some generalised cohomology theories were obtained in [17]. Motivated by these results, one can conjecture:
@@
Let $L\subseteq M$ be a compact, relatively exact Lagrangian in a (compact or Liouville) symplectic manifold $M$.
Let $\tilde{\phi}\mathrel{\mathop{\ordinarycolon}}L\to L$ be a diffeomorphism which extends to a Hamiltonian diffeomorphism $\phi\mathrel{\mathop{\ordinarycolon}}M\to M$. Then $\tilde{\phi}$ is isotopic to the identity.
Remark 1.5.
Theorem 1.3 only constrains the homotopy class of $\tilde{\phi}$. Using different methods (namely generating functions), upcoming work with Courte [3] finds constraints on the smooth isotopy class of $\tilde{\phi}$ in the setting of Theorem 1.3 not detectable by their homotopy class, in the special case $L$ is the zero-section in the cotangent bundle of a high-dimensional torus $M=T^{*}T^{n}$.
1.2 $C^{0}$ symplectic topology
@@
A symplectic diffeomorphism $\theta\mathrel{\mathop{\ordinarycolon}}M\to M$ is a diffeomorphism that preserves the symplectic structure, meaning $\theta^{*}\omega=\omega$.
Any Hamiltonian diffeomorphism is a symplectomorphism [16, Proposition 10.2].
$C^{0}$ symplectic topology studies non-smooth generalisations of Hamiltonian and symplectic diffeomorphisms:
Definition 1.7.
A symplectic homeomorphism is a homeomorphism $\theta\mathrel{\mathop{\ordinarycolon}}M\to M$ which is a $C^{0}$ limit of symplectic diffeomorphisms.
A Hamiltonian homeomorphism is a homeomorphism $\theta\mathrel{\mathop{\ordinarycolon}}M\to M$ which is a $C^{0}$ limit of Hamiltonian diffeomorphisms. We write $\overline{Ham}(M)$ for the space of Hamiltonian homeomorphisms on $M$.
The field of $C^{0}$ symplectic topology began with the following theorem.
Theorem 1.8 (Eliashberg-Gromov [4, 7]).
Any symplectic homeomorphism which is also a diffeomorphism is in fact a symplectic diffeomorphism.
This is surprising since the condition of being a symplectic diffeomorphism involves the derivative of $\theta$, and in general one often loses control of derivatives under $C^{0}$ limits. The Hamiltonian version of this theorem, namely that any Hamiltonian homoeomorphism which is a diffeomorphism is a Hamiltonian diffeomorphism, remains open.
In general, the geometry of Hamiltonian homeomorphisms is much less understood than those of Hamiltonian diffeomorphisms; as we illustrate in the following examples, $C^{0}$ symplectic topology both exhibits in some ways similar rigidity phenomena to smooth symplectic topology, and in other ways much more flexibility. It is this interplay which makes it such an intriguing subject.
Example 1.9 (Rigidity in $C^{0}$ symplectic topology).
Seidel and Keating [21, 13] produced large classes of examples of symplectomorphisms which are smoothly isotopic to the identity, but not isotopic to the identity through (smooth) symplectomorphisms. Roughly, these are constructed as compositions of symplectic Dehn twists. Jannaud [12] later showed that many of these are also not isotopic to the identity through symplectic homeomorphisms.
Example 1.10 (Flexibility in $C^{0}$ symplectic topology).
One of the main goals in (smooth) symplectic topology has been to find lower bounds on the numbers of fixed-points of Hamiltonian diffeomorphisms $f\mathrel{\mathop{\ordinarycolon}}M\to M$. Motivated by the Arnol’d conjecture, it has been shown [6, 14, 19] that if $M$ is compact and $f$ is generic, there must be at least as many fixed-points as the sum of the Betti numbers of $M$; in particular, there must be at least 2.
It is natural to ask whether the same inequality holds for Hamiltonian homeomorphisms. Surprisingly, this fails badly: Buhovsky-Humilière-Seyfaddini [2] showed that whenever $\mathrm{Dim}(M)\geq 4$, there exists a Hamiltonian homeomorphism $\theta\mathrel{\mathop{\ordinarycolon}}M\to M$ with a single fixed point.
1.3 Monodromy in the $C^{0}$ setting
Here we extend parts (1) and (2) of Theorem 1.3 to the $C^{0}$ setting (it would be interesting to extend parts (3) and (4) to the $C^{0}$ setting too).
@@ 1.3 Monodromy in the $C^{0}$ setting
Let $L\subseteq M$ be a compact, relatively exact Lagrangian in a (compact or Liouville) symplectic manifold $M$.
Let $\tilde{\phi}\mathrel{\mathop{\ordinarycolon}}L\to L$ be a homeomorphism, and assume that there is a Hamiltonian homeomorphism $\phi\mathrel{\mathop{\ordinarycolon}}M\to M$ extending $\tilde{\phi}$.
@@
Theorem 1.11 holds for various other generalised cohomology theories $R^{*}$ (such as real $K$-theory $KO^{*}$) provided the condition [17, Proposition 1.14] holds for $R^{*}$. Examples of diffeomorphisms of closed manifolds which act as the identity but not on (real or complex) $K$-theory are given in [17, Appendices A & B].
@@
The Nearby Lagrangian conjecture says that any closed exact Lagrangian $L$ in a cotangent bundle $T^{*}Q$ is Hamiltonian isotopic to the zero-section $Q_{0}\subseteq T^{*}Q$. As we show in Section 2, a positive resolution to this conjecture would allow us to deduce Theorem 1.11 from Theorem 1.3.
Remark 1.13.
Question 1.1 can be viewed as the study of the relative mapping class group $\pi_{0}\mathrm{Ham}(M,L)$ of isotopy classes of Hamiltonian diffeomorphisms which fix $L$, and Theorem 1.11 is then the corresponding extension to Hamiltonian homeomorphisms $\pi_{0}\overline{\mathrm{Ham}}(M,L)$.
In a slightly different direction, homotopy clases of loops of Hamiltonian diffeomorphisms (equivalently, nontrivial elements of $\pi_{1}\mathrm{Ham}(M)$) were constructed in [20] and shown to be nontrivial using the Seidel representation. This has been extended recently in [11] to show that these loops are still nontrivial as homotopy classes of Hamiltonian homeomorphisms (equivalently, nonzero in $\pi_{1}\overline{Ham}(M)$), by combining the Seidel representation with spectral invariants.
@@ 1.4 Extension to the relative setting
Assume $\tilde{\phi}\mathrel{\mathop{\ordinarycolon}}L\to L$ is a homeomorphism which extends to a Hamiltonian homeomorphism $\phi\mathrel{\mathop{\ordinarycolon}}M\to M$. $\phi$ induces a map of pairs $(M,L)\to(M,L)$, and hence a map on relative cohomology.
Conjecture 1.4 would imply that $\phi$ is homotopic to the identity as a map of pairs; we do not prove this, but prove the cohomological analogue:
@@
Let $L\subseteq M$ be a compact, relatively exact Lagrangian in a (compact or Liouville) symplectic manifold $M$.
Let $\tilde{\phi}\mathrel{\mathop{\ordinarycolon}}L\to L$ be a homeomorphism, and assume that there is a Hamiltonian homeomorphism $\phi\mathrel{\mathop{\ordinarycolon}}M\to M$ extending $\tilde{\phi}$. Then:
@@
Remark 1.15.
It is tempting to attempt to deduce Theorem 1.14 the long exact sequence on cohomology associated to the pair $(M,L)$, and use Theorem 1.11 along with the fact $\phi$ must act as the identity on $H^{*}(M)$.
However, this does not follow automatically, because the five lemma does not constrain automorphisms.
[…]
We illustrate a concrete example of this phenomenon, in the following commutative diagram:
[…]
The proof of Theorem 1.14 is similar to the proof of Theorem 1.11, but whereas the proof of Theorem 1.11 uses a moduli space of holomorphic curves with domain $Z_{C}$, $\mathcal{P}_{C}$, along with the evaluation map on the boundary of $Z_{C}$, $ev\mathrel{\mathop{\ordinarycolon}}\mathcal{P}\times\partial Z_{C}\to M$, the proof of Theorem 1.14 uses the evaluation map on the whole of $Z_{C}$, $ev\mathrel{\mathop{\ordinarycolon}}\mathcal{P}\times Z_{C}\to M$.
@@ 1.5 Acknowledgements
The author thanks Amanda Hirschi, Ivan Smith and Yuhan Sun for helpful conversations, Jack Smith for helpful conversation and useful comments on an earlier draft, and the anonymous referee for useful comments and feedback. The author is supported by the Engineering and Physical Sciences Research Council [EP/W015889/1].
2 Relationship to the Nearby Lagrangian conjecture
One of the main open problems in symplectic topology is:
Conjecture 2.1 (Nearby Lagrangian Conjecture).
Any closed exact Lagrangian $L$ in a cotangent bundle $T^{*}Q$ is Hamiltonian isotopic to the zero-section $Q_{0}\subseteq T^{*}Q$.
As we show in Proposition 2.2 below, if the Nearby Lagrangian conjecture holds, we can reduce the exact case of Theorem 1.11 to Theorem 1.3.
Proposition 2.2.
Suppose $(M,\omega)$ is a Liouville symplectic manifold and $L$ is a closed exact Lagrangian in $M$. Let $\tilde{\phi}\mathrel{\mathop{\ordinarycolon}}L\to L$ be a homeomorphism that extends to a Hamiltonian homeomorphism $\phi\mathrel{\mathop{\ordinarycolon}}M\to M$.
Assume the Nearby Lagrangian Conjecture holds. Then $\tilde{\phi}$ is homotopic to a diffeomorphism $\tilde{\psi}\mathrel{\mathop{\ordinarycolon}}L\to L$ that extends to a Hamiltonian diffeomorphism $\psi\mathrel{\mathop{\ordinarycolon}}M\to M$.
Using Proposition 2.2, one may then apply results such as Theorem 1.11, which only apply to Hamiltonian diffeomorphisms, to $\tilde{\psi}$ to deduce conclusions about the homotopy class of $\tilde{\phi}$.
Proof of Proposition 2.2.
Weinstein’s neighbourhood theorem [16, Theorem 3.33] says there is a tubular neighbourhood $U\subseteq M$ of $L$ which is symplectomorphic to the disc bundle $D^{*}_{r}L$ of some radius $r$. Let $\pi\mathrel{\mathop{\ordinarycolon}}U\to L$ be the projection map.
By definition, we may choose a Hamiltonian diffeomorphism $\phi^{\prime}\mathrel{\mathop{\ordinarycolon}}M\to M$ which is $C^{0}$-close to $\phi$. $\phi^{\prime}$ may not send $L$ to itself, but (for $\phi^{\prime}$ sufficiently $C^{0}$-close to $\phi$), $\phi^{\prime}(L)$ does lie in $U$, and $\pi\circ\phi^{\prime}\mathrel{\mathop{\ordinarycolon}}L\to L$ is $C^{0}$-close to (and hence homotopic to) $\phi$. By the Nearby Lagrangian conjecture, we may assume $L^{\prime}$ is Hamiltonian isotopic to $L$ in $U$. By isotopy extension for Hamiltonians (similarly to [16, Exercise 3.40]), there is a Hamiltonian isotopy from $L^{\prime}$ to $L$ in $M$, staying entirely in the Weinstein neighbourhood $U$. Let $\theta$ be the corresponding Hamiltonian diffeomorphism of $M$, sending $L^{\prime}$ to $L$. Let $\psi=\theta\circ\phi^{\prime}$; this sends $L$ to $L$, and is a Hamiltonian diffeomorphism by [16, Proposition 10.2]. We set $\tilde{\psi}=\psi|_{L}\mathrel{\mathop{\ordinarycolon}}L\to L$. It remains to show $\tilde{\psi}$ is homotopic to $\tilde{\phi}$.
Since $\phi^{\prime}$ is $C^{0}$-close to $\phi$, it follows that $\pi\circ\phi^{\prime}\mathrel{\mathop{\ordinarycolon}}L\to L$ is $C^{0}$-close to, and hence homotopic to, $\tilde{\phi}$.
Since the Hamiltonian isotopy from $L^{\prime}$ to $L$ lies in $U$, by composing with $\pi$ we obtain a homotopy from $\pi\circ\phi^{\prime}$ to $\tilde{\psi}$.
Combining these final two homotopies, we obtain one between $\tilde{\psi}$ and $\tilde{\phi}$. ∎
Remark 2.3.
Let $M=T^{*}Q$ be a cotangent bundle. The strong nearby Lagrangian conjecture says that the space of closed exact Lagrangians in $M$ is contractible. This would imply Conjecture 1.4 in the case $L\subseteq M$ is the zero-section.
@@ 3 Smooth approximation
We first choose a complete Riemannian metric on $M$, and let $2\varepsilon>0$ be its injectivity radius. Write $d$ for the Riemannian distance function. We may assume that $L$ is a totally geodesic submanifold with respect to this metric.
Now choose a smooth Hamiltonian diffeomorphism $\psi\mathrel{\mathop{\ordinarycolon}}M\to M$ which is sufficiently $C^{0}$-close to $\phi$ that for all $x$ in $M$, $d(\psi(x),\phi(x))\leq\varepsilon$. We write $L^{\prime}$ for $\psi(L)$.
Let $U$ be a tubular neighbourhood of $L$ and $\pi\mathrel{\mathop{\ordinarycolon}}U\to L$ the projection map. We choose $U$ small enough that for all $x$ in $U$,
@@
We may choose $\psi$ $C^{0}$-close enough to $\phi$ such that $L^{\prime}$ lives in $U$, and generically so that $L$ and $L^{\prime}$ intersect transversally.
@@
The maps $\phi,\pi\circ\psi\mathrel{\mathop{\ordinarycolon}}L\to L$ are homotopic.
@@
For $x$ in $L$, since $\pi\circ\phi(x)=\phi(x)$, by the triangle inequality $d(\phi(x),\pi\circ\psi(x))\leq 2\varepsilon$. Thus there is a unique small geodesic between $\phi(x)$ and $\pi\circ\psi(x)$; since $L\subseteq M$ is totally geodesic, this geodesic stays in $L$. Moving along this geodesic for each $x$ gives the required homotopy. ∎